OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment whose endpoints are H(8, 13) and K(10, 9).

OpenStudy (astrophysics):

\[\large \left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\right)\] Midpoint formula

OpenStudy (anonymous):

thank you